OpenStudy (anonymous):

I'm trying to proof that c needs to be equal to 4 for the following equality. n is all positive integers But I'm getting 2, can anyone pin point my mistake? \[\left| n^4-3n^6\right|<= cn^6\] \[-n^4 + 3n^6 <= cn^6\] \[-1+3n^2<=cn^2\] \[-1 <= n^2(c-3)\] \[\sqrt{-1/(c-3)} <= n\]

OpenStudy (anonymous):

where did the negative from the absolute to n^4 come from

OpenStudy (anonymous):

Since n represents only positive numbers then I know that, \[\left| n^4-3n^6 \right| = -n^4+3n^6\]

OpenStudy (anonymous):

since n^4 and 3n^4 will always be positive are you trying to say?

OpenStudy (anonymous):

The result yes and since n^6 is of a higher degree I'm setting it to be positive. For example, if n = 1 then Abs(1-3) = 2 which is the same as -1 + 3

OpenStudy (anonymous):

Now that I look at it, I think c should be 3

OpenStudy (ybarrap):

I get 2 as well: ∣n4−3n6∣∣<=cn6, note that c >= 0 -cn6 <= n4 - 3n6 <= cn6 -cn6 + 3n6 <= n4 <= cn6 + 3n6 (3-c)n6 <= n4 <= (3+c)n6 (3-c)n2 <= 1 <= (3 +c)n2 1<= (3 + c)n2 and (3-c)n2 <= 1 1/n2 - 3 <= c and 1/n2 - 3 >= -c c >= 1/n2 - 3 and c >= 3 - 1/n2 => c >= 3 - 1/n2 >= 3 - 1 = 2, since smallest n = 1 Smallest possible c is 2 largest possible c is lim n -> inf 3 - 1/n2 = 3, but not quite reaching 3 so largest integer c can be is 2 Hence c = 2.

OpenStudy (anonymous):

Though call. It clearly cannot be 2, but maybe for a large n it is but I just plugged n = 10000000000 and it wasn't even close. Using the limit I can proof c = 3. Thanks.

OpenStudy (ybarrap):

I should have taken as supremum 3 in my previous attempt. Here is another look. ∣n4−3n6∣ <= cn6, note that c >= 0 because |n4 - 3n6| >= 0, then cn6 >= 0 For n = 1 (smallest possible value) |n4 - 3n6| = |1 - 3| = 2 But then this implies |n4 - 3n6| = 2 <= cn6 = c and c >= 2. c = 1 is not possible. For n=2: |n4 -3n6| = |16 - 3*64| = 16|1 - 3*4| = 16(11) < = c2^6 = 6 so c >= 16*11/64 = 11/4 ==> c >= 3, since c must be an integer As n --> inf 0 <= |n4 - 3n6|/n6 <= c 0 <= |0 - 3| <= c 0 <= 3 <= c Again, c >=3. Hence least upper bound on c is 3.

OpenStudy (anonymous):

This is what i got solving for c also, ... however i didn't change the signs because you end up with an imaginary number